Rewrite 2/9 plus 5/12 with a common denominator

Answer by Amygdala(1)

2/9 + 5/12 = 2(12)/9(12) (Multiply both the reciprocal and denominator by 12)
+ 5(9)/12(9) (Multiply this reciprocal and denominator by 9 so we can get same common denominator)
= 24/108 + 45/108 = 69/108